Which of the following coordinates would be included in the graph

y = 3x + 2?
(There may be more than one correct answer.)


A. (0, 0)
B. (2, 4)
C. (-1,-1)
D. (6, 3)

Answer by nerdybill(7384)

y = 3x + 2
The only solution that works is:
C. (-1,-1)
.
y = 3x + 2
-1 = 3(-1) + 2
-1 = -3 + 2
-1 = -1